Car A has an acceleration of 2 m/s^2 due east and car B, 4 m/s^2 due north. What is the acceleration of car B with respect to car A?

Grade:11

3 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Chanchal

aA=2i

aB=4j

rel acceleration = aB - aA= 4j-2i

magnitude=√42+22=√20 m/s2

direction=tan-1(4/2)=tan-12=63.430 from west towards north

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

vikas askiitian expert
13 years ago

accleration of car A is a = 2i              (due to east)

accleration of car B is b =4j              (due to north)

accleration of B wrt A is arel = a-b

 arel = 2i - 4j

 a = (22+42)1/2 = 2(sqrt5)m/s2 in magnitude

Deepak Sharma
13 years ago
accleration of car A is a = 2i (due to east) accleration of car B is b =4j (due to north) accleration of B wrt A is arel = a-b arel = 2i - 4j a = (22+42)1/2 = 2(sqrt5)m/s2 in magnitude